Please Help me I never was good at word problem. Please show all work.
Thanks in advance
The length of a rectangular garden is 7 feet longer than its width. If the garden's perimeter is 202 feet, what is the area of the garden in square feet?

Answer by Alan3354(69443)

The length of a rectangular garden is 7 feet longer than its width. If the garden's perimeter is 202 feet, what is the area of the garden in square feet?
-------
Find L & W
---
Perimeter = 2L + 2W = 202
L + W = 101
L = W + 7
------
Now it's not a word problem.